OpenStudy (anonymous):

consider the following quadratic equation: x^2-16x+64=0 using the standard form ax^2+bx+c=0 of the quadratic equation above, factor the left hand side of the equation into two linear factors.

OpenStudy (anonymous):

(x-8)^2 = 0

OpenStudy (anonymous):

recognize a "perfect square" as \[a^2+2ab+b^2=(a+b)^2\] in this example \[a=x,b=-8\]